# Scaling Instruction-Based Video Editing with a High-Quality Synthetic Dataset

### v1

@@ 4 Model Training via Modality Curriculum Learning
We select the in-context video generator VACE (Jiang et al., 2025) as our backbone, inspired by its strong prior for generating videos that are spatially and structurally aligned with a source video. VACE’s original capability is to condition its generation on two visual contexts (and prompts): a source video and a reference image. Our goal is to repurpose this powerful visual generator into a proficient editor that operates on abstract textual instructions. However, directly fine-tuning the model to bridge the vast semantic gap from visual to textual conditioning is prone to instability. We therefore adapt its architecture, as shown in Fig. 4. It consists of a Context Branch for extracting spatiotemporal features from the source video and reference frame, and a DiT-based (Peebles & Xie, 2023) Main Branch that synthesizes the edited video under the joint guidance of the visual context and the new textual embeddings from the instruction.
@@
To ease the training difficulty and stably bridge this modality gap, we introduce a modality curriculum learning (MCL) strategy. The core idea is to leverage the model’s inherent ability to process the reference image context as a temporary aid. In the initial training phase, we provide the edited reference frame as a strong visual ”scaffold” alongside the new text instruction. As training progresses, we gradually anneal the probability of providing this visual scaffold, eventually dropping it entirely. This process compels the model to shift its dependency from the concrete visual target it already understands to the more abstract textual instruction, transforming it into a purely instruction-based video editing model. We train the model using the flow matching (Lipman et al., 2022) objective:
@@
Figure 5: Qualitative comparisons with prior arts TokenFlow (Geyer et al., 2024), InsV2V (Cheng et al., 2024), InsViE (Wu et al., 2025b) and Gen4-Aleph (Runway, 2025).
@@ 5 Experiments
5.1 Experimental Settings
Our model is built upon the pre-trained in-context video generator (Jiang et al., 2025; Wan et al., 2025) and is fine-tuned on our newly proposed large-scale dataset, which comprises over one million high-quality video triplets. To maintain the strong generative prior of the base model and ensure training efficiency, we freeze the majority of the pre-trained model’s parameters, and only fine-tune the linear projection layers of context blocks. The model is trained for approximately 16,000 steps using the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of 1e-4 on a cluster of 64 NVIDIA H-series GPUs. We employ our modality curriculum learning strategy, where the initial 5,000 steps serve as a curriculum warm-up phase.
@@ 5.2 Experimental Results
Quantitative Comparison. We perform quantitative comparisons using automatic metrics and a user study, summarized in Table 2. For automatic evaluation, we employ three metrics: CLIP-T measures the CLIP text-video similarity to assess how well the edit follows the instruction; CLIP-F calculates the average inter-frame CLIP similarity to gauge temporal consistency; and a VLM score provides a holistic assessment of edit effectiveness, semantic preservation, and overall aesthetic quality. A user study based on 1,000 ratings also rates instruction-following (Edit-Acc), temporal consistency (Temp-Con), and overall quality (Overall). As shown, our method significantly outperforms all baselines across metrics, achieving the highest automatic scores and a strong preference in human evaluations, which confirms its superior instruction adherence, temporal smoothness, and visual quality. Please refer to Appendix C for additional details of the user study.
@@
Additional Results. We showcase the synthetic-to-real (sim2real) capability in Fig. 6 benefited from our data by training the model to map the stylized videos in our dataset back to their original, real-world source videos. This successful transfer highlights the rich and photorealistic information contained within our dataset, demonstrating its utility beyond standard editing tasks. Also, our final trained model substantially outperforms the raw data generator from our pipeline, demonstrating superior handling of newly emerged content as in Fig. 7. This superiority stems from our scaled training regimen, including the curriculum learning and exposure to the filtered, high-quality data.
@@ 6 Conclusion
We have presented Ditto, a scalable framework that significantly advances instruction-based video editing by systematically addressing the core challenge of data scarcity through a new paradigm for large-scale data synthesis. Our synthetic data generation pipeline overcomes the fidelity-diversity and efficiency-coherence trade-offs plaguing prior methods by leveraging strong image-editing priors, a distilled in-context video generator with a temporal enhancer, and autonomous VLM-based quality control. This enables the creation of the large-scale, high-quality Ditto-1M dataset. The proposed modality curriculum learning strategy further ensures our model Editto achieves state-of-the-art performance by effectively transitioning from visual-textual conditioning to purely instruction-driven inference. Extensive results reveal the effectiveness of the proposed method.
Appendix A Appendix Overview
In this appendix, we provide additional details and results to supplement our main paper. In Appendix B, we present an analysis of the in-context data generator. We showcase the interface used for our human evaluations in Appendix C. Finally, Appendix D contains more qualitative results of our dataset and model, and a word cloud visualizing the distribution of instructions in our dataset.
Appendix B Justifying the Design of Data Generation Pipeline
Figure 9: Results of various settings for data generation.
We provide an analysis of the videos synthesized by the in-context video generator, VACE, to justify the design of our data pipeline. As in Fig. 9, we first observe that using only depth maps to guide the generator results in a significant loss of content from the source video, leading to poor fidelity. Conversely, conditioning the generator on a keyframe from the original source video alongside the editing instruction fails to produce the desired edit - the output remains almost identical to the source. These findings reveal that while the base generator excels at motion transfer, its inherent instruction-following capability for editing tasks is limited. Based on this analysis, we validate our proposed approach: using a keyframe modified by an advanced image editor, in conjunction with depth guidance as the context. This method achieves the optimal balance of instruction adherence, temporal consistency, and source fidelity for our data synthesis.
Appendix C Demonstration of User Study Interface
Figure 10: The interface of the user study.
To collect human preference data, we designed a user-friendly evaluation interface, as shown in Fig. 10. For each source video and text prompt, we presented participants with the edited results from different methods in a randomized order. They were then asked to rank the videos from best (1) to worst (4) based on three criteria: Instruction Following, Temporal Consistency, and Overall preference. The final scores are calculated based on the ranking.
Figure 11: Additional visualization of data from the proposed dataset and model outputs. Please view the site for additional video samples.
Appendix D Additional Results of Dataset and Model
Figure 12: The word cloud of editing instructions.
We present additional qualitative results to further demonstrate our dataset and model’s performance across a wide range of editing instructions in Fig. 11. We also include a word cloud in Fig. 12 that illustrates the diversity and distribution of the editing prompts within the dataset, highlighting its comprehensive coverage.

### v2

@@ 4 Modality Curriculum Model Learning
We select the in-context video generator VACE [jiang2025vace] as our backbone, inspired by its strong prior for generating videos that are spatially and structurally aligned with a source video. VACE’s original capability is to condition its generation on two visual contexts (and prompts): a source video and a reference image. Our goal is to repurpose this powerful visual generator into a proficient editor that operates on abstract textual instructions. However, directly fine-tuning the model to bridge the vast semantic gap from visual to textual conditioning is prone to instability. We therefore adapt its architecture, as shown in Fig. 3. It consists of a Context Branch for extracting spatiotemporal features from the source video and reference frame, and a DiT-based [peebles2023dit] Main Branch that synthesizes the edited video under the joint guidance of the visual context and the new textual embeddings from the instruction.
@@
To ease the training difficulty and stably bridge this modality gap, we introduce a modality curriculum learning (MCL) strategy. The core idea is to leverage the model’s inherent ability to process the reference image context as a temporary aid. In the initial training phase, we provide the edited reference frame as a strong visual ”scaffold” alongside the new text instruction. As training progresses, we gradually anneal the probability of providing this visual scaffold, eventually dropping it entirely. This process compels the model to shift its dependency from the concrete visual target it already understands to the more abstract textual instruction, transforming it into a purely instruction-based video editing model. We train the model using the flow matching [lipman2022flowmatch] objective:
@@ 5 Experiments
5.1 Experimental Settings
Our model is built upon the pre-trained in-context video generator [jiang2025vace, wan2025] and is fine-tuned on our newly proposed large-scale dataset, which comprises over one million high-quality video triplets. To maintain the strong generative prior of the base model and ensure training efficiency, we freeze the majority of the pre-trained model’s parameters, and only fine-tune the linear projection layers of context blocks. The model is trained for approximately 16,000 steps using the AdamW optimizer [loshchilov2017adamw] with a constant learning rate of 1e-4 on a cluster of 64 GPUs. We employ our modality curriculum learning strategy, where the initial 5,000 steps serve as a curriculum warm-up phase.
@@ 5.2 Experimental Results
Quantitative Comparison. We perform quantitative comparisons using automatic metrics and a user study, summarized in Tab. 1. To ensure a fair comparison, our test set consists of 50 videos collected from various online sources, deliberately excluding Pexels videos to ensure the data is out-of-distribution relative to our training set. For each video, we provide 5 distinct editing instructions. For automatic evaluation, we employ three metrics: CLIP-T measures the CLIP text-video similarity to assess how well the edit follows the instruction; CLIP-F calculates the average inter-frame CLIP similarity to gauge temporal consistency; and a VLM score provides a holistic assessment of edit effectiveness, semantic preservation, and overall aesthetic quality. A user study based on 1,000 votes from postgraduates and researchers also rates instruction-following (Edit-Acc), temporal consistency (Temp-Con), and overall quality (Overall). As shown, our method significantly outperforms all baselines across metrics, achieving the highest automatic scores and a strong preference in human evaluations, which confirms its superior instruction adherence, temporal smoothness, and visual quality. Please refer to the supplementary materials for details of the user study.
@@
Additional Results. We showcase the synthetic-to-real (syn2real) capability in Fig. 5 benefited from our data by training the model to map the stylized videos in our dataset back to their original, real-world source videos. This successful transfer highlights the rich and photorealistic information contained within our dataset, demonstrating its utility beyond standard editing tasks. Also, our final trained model substantially outperforms the raw data generator from our pipeline, demonstrating superior handling of newly emerged content as in Fig. 6. This superiority stems from our scaled training regimen, including the curriculum learning and exposure to the filtered, high-quality data.
@@ 6 Conclusion
We have presented Ditto, a scalable framework that significantly advances instruction-based video editing by systematically addressing the core challenge of data scarcity through a new paradigm for large-scale data synthesis. Our synthetic data generation pipeline overcomes the fidelity-diversity and efficiency-coherence trade-offs plaguing prior methods by leveraging strong image-editing priors, a distilled in-context video generator with a temporal enhancer, and autonomous VLM-based quality control. This enables the creation of the large-scale, high-quality Ditto-1M dataset. The proposed modality curriculum learning strategy further ensures our model Editto achieves state-of-the-art performance by effectively transitioning from visual-textual conditioning to purely instruction-driven inference.